OpenStudy (anonymous):

Angles C and D are contemporary angles. Determine the measures of Angles C and D if angle D is 15 degrees less than twice angle C.

Directrix (directrix):

Let the measure of angle C = x. Then, <D = 2x - 15. Because the angles are complementary, x + 2x - 15 = 90 Solve for x.

OpenStudy (anonymous):

would it be 2x^2-15=90 +15 +15 --------------- 2x^2= 105 ?..

OpenStudy (anonymous):

or.. 3x?

OpenStudy (anonymous):

3x=105 --- ---- 3. 3 x= 35 ?

Directrix (directrix):

> x= 35 ? This is correct. Now, we have to get the angle measures.

Directrix (directrix):

angle C = x = ? D is 15 degrees less than twice angle C = 2x - 15 = ?

OpenStudy (anonymous):

C= 2 (35)'15?

OpenStudy (anonymous):

-15*

OpenStudy (anonymous):

C= 55?

OpenStudy (anonymous):

is 35 angle C? 0.0

Directrix (directrix):

angle C = x = 35. D is 15 degrees less than twice angle C = 2*35 - 15 = 55

Directrix (directrix):

> is 35 angle C? YES

OpenStudy (anonymous):

thank you for your help.